OpenStudy (anonymous):

Joseph wants to estimate 1 3/7 * 5 4/3 To estimate, he simplifies the expression 1 1/2 * 6

OpenStudy (anonymous):

seems like it just wants you to solve \[1\frac{ 1 }{ 2 }*6\]

OpenStudy (anonymous):

this can be written as \[\frac{ 3 }{ 2 }*6\]

OpenStudy (anonymous):

Opps I forgot to add the rest >.< Is Joseph’s estimate a close estimate, an underestimate, or an overestimate? A. underestimate B. overestimate C. close estimate

OpenStudy (anonymous):

;-;

OpenStudy (anonymous):

ok then you'll want to figure out the answers to\[1\frac{ 3 }{ 7 }*5\frac{ 4 }{ 3 }\] and\[1\frac{ 1 }{ 2 }*6\]

OpenStudy (anonymous):

then compare the two answers. changing mixed fractions into normal fractions can be done by multiplying the number in front of fraction with the denominator and adding it with the numerator and then put that quantity over the denominator. For example \[1\frac{ 3 }{ 7 }=\frac{ 7*1+3 }{ 7 }=\frac{ 10 }{ 7 }\]

OpenStudy (anonymous):

I got 19/3 for 5 4/3

OpenStudy (anonymous):

\[10/7 * 19/3 = 190/21\]

OpenStudy (anonymous):

correct now do the same for the other equation and compare the two. It might be nice to change your answer into decimal format using a calculator to make comparison easier.

OpenStudy (anonymous):

Okay thank you. :)

OpenStudy (anonymous):

9.0476190 and 1.5 so the answer is A.?

OpenStudy (anonymous):

Hallo?

OpenStudy (anonymous):

yes thats looks correct

OpenStudy (anonymous):

well done @Obbie